1/2

Edward Lear,  Edward Lear,
The Owl and the Pussy-cat  La Strig’ kaj la Kato

translated by Pejno Simono

I  I

The Owl and the Pussy-cat went to sea In a beautiful pea-green boat, They took some honey, and plenty of money, Wrapped up in a five pound note. The Owl looked up to the stars above, And sang to a small guitar, “O lovely Pussy! O Pussy my love, What a beautiful Pussy you are, You are, You are! What a beautiful Pussy you are!”

La Strig’ kaj la Kato en verda boato ekvelis kun kargo abunda, mielo por manĝi kaj mono por ŝanĝi en Angla bileto kvinpunda. Sub nokta stelaro kun eta gitaro ekkantis la Strig’ pro inklin’, “Ho bela vi Kato! Ho mia amato! Ho Kato vi ravas ja min, ja min, ja min! Ho Kato vi ravas ja min!”

II  II

Pussy said to the Owl, “You elegant fowl! How charmingly sweet you sing! O let us be married! too long we have tarried: But what shall we do for a ring?” They sailed away, for a year and a day, To the land where the Bong-tree grows And there in a wood a Piggy-wig stood With a ring at the end of his nose, His nose, His nose, With a ring at the end of his nose.

La Kat’ aklamanta “Vi bird’ eleganta! Vi kantas ĉarmege ulule! Ho ni geedziĝu! Hezit’ ne estiĝu! Sed havas da ringoj ni nule!” Tra jaro plus tago ekestis vojaĝo al land’, kie kreskas bonguj’, kaj jen en arbaro renkontis la paro Porkidon kun ring’ en naztru’, naztru’, naztru’. Porkidon kun ring’ en naztru’.

…  …

2/2

III  III

“Dear Pig, are you willing to sell for one shilling Your ring?” Said the Piggy, “I will.” So they took it away, and were married next day By the Turkey who lives on the hill. They dined on mince, and slices of quince, Which they ate with a runcible spoon; And hand in hand, on the edge of the sand, They danced by the light of the moon, The moon, The moon, They danced by the light of the moon.

“Jen unu ŝilingo por via nazringo! Ĉu bone?” La Porko kvitigis. La tagon pli poste la paron senkoste Kok’ Hinda sur monto ĵurligis. De ment’ la foliojn kaj cent cidoniojn ili manĝis per runca kuler’, festenis lunlume kaj dancis brakume sur plaĝ’ en la nokta aer’, aer’, aer’. Sur plaĝ’ en la nokta aer’.

Author of this English poem is Edward Lear.

Arg-976-1966 (2014-03-13 21:55:44)

Pri la verkintgo vidu la retejon http: // de. wikipedia. org/ wiki/ Edward_Lear respektive http: // eo. wikipedia. org/ wiki/ Edward_Lear .

Translation of the English poem “The Owl and the Pussy-cat” by Edward Lear into Esperanto by Pejno Simono (Simon Edward Adrian Payne, ∗1941-12-23).

Arg-976-1967 (2014-03-13 21:45:19)

Tiu ĉi poem-esperantigo troveblas en la retejo http: // www. bompa. org/ .